34) What are the units of k in the following rate law? 

Rate =k[x]2[Y] 


35) What are the units of k in the following rate law? 

Rate = k[X] [Y]3


36) What are the units of k in the following rate law? 

Rate = k[X][Y] 


37) What are the units of k in the following rate law? 

Rate = k[X] 


38) What are the units of k in the following rate law? 

Rate = k[X]0[Y]0
Answer #1

34)
Rate = k[X]^2[Y]
M.s-1 = k (M)^2*(M)
M.s-1 = k (M)^3
k = M-2.s-1
Answer: M-2.s-1

35)
Rate = k[X][Y]^3
M.s-1 = k (M)*(M)^3
M.s-1 = k (M)^4
k = M-3.s-1
Answer: M-3.s-1

36)
Rate = k[X][Y]
M.s-1 = k (M)*(M)
M.s-1 = k (M)^2
k = M-1.s-1
Answer: M-1.s-1

37)
Rate = k[X]
M.s-1 = k (M)
k = s-1
Answer: s-1

38)
Rate = k[X]o[Y]o
Rate = k
M.s-1 = k
k = M.s-1
Answer: M.s-1
Answer #2

putang ina mo maramot mukhang pera